# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 2.1.4   Electroproduction
Meson photoproduction has become an essential tool in the search for new excited light-quark baryon states. As discussed in the previous section, many new excited states have been discovered thanks to high-precision photoproduction data in different final states [283], and are now included in recent editions of the Review of Particle Physics (RPP) [284]. The experimental exploration of resonance transition form factors spans over 60 years, with many review articles describing this history. Here we refer to some recent ones [15, 285, 9, 286]. A review of recent electroproduction experiments in hadron physics and their interpretation within modern approaches to strong interaction physics can be found in Ref. [287]. The exploration of the internal structure of excited states and the effective degrees of freedom contributing to $s$-channel resonance excitation requires the use of electron beams, where the virtuality ($Q^{2}$) of the exchanged photon can be varied to penetrate through the peripheral meson cloud and probe the quark core and its spatial structure. Electroproduction can thus reveal whether a resonance is generated through short-distance photon interaction with the small quark core or through interaction with a more extended hadronic system. Electroproduction of final states with pseudoscalar mesons (e.g., $N\pi$, $p\eta$, $K\Lambda$, $p\pi^{-}\pi^{+}$) has been employed at Jefferson Laboratory mainly with the CEBAF Large Acceptance Spectrometer (CLAS) operating at an instantaneous luminosity of $10^{34}\,\mathrm{sec}^{-1}\mathrm{cm}^{-2}$. In Hall A and Hall C, pairs of individual well-shielded focusing magnetic spectrometers are employed with more specialized aims and limited acceptance, but operating at much higher luminosity. This experimental program has led to new insights into the scale dependence of effective degrees of freedom, e.g., meson-baryon, constituent quarks, and dressed quark contributions. Data on meson electroproduction can be found in Refs. [288, 289, 290, 291, 292, 152, 146, 293, 294, 295, 296, 297, 298, 299, 300, 301, 302, 303, 304, 305, 306, 307, 308, 309, 310, 311, 312, 313, 314, 315, 316, 286, 317, 318, 319, 15, 320, 321, 322, 323, 324, 325, 326, 327, 328, 329, 330, 331, 332, 210, 333, 334, 335, 336, 337, 338, 9, 339, 218, 340, 341, 342, 343, 344, 345, 346, 347, 348, 349, 350, 351, 352, 353, 354].
@@ 2.2   Analyses
2.2.1   Partial-wave analyses and coupled-channel approaches
@@
The main KH results can be found in [21], with an update of the results presented in [355]. A comprehensive description of methods and results can be found in the handbook [24] published in the Landold-Börnstein series. Höhler presented updated resonance masses and widths using the so-called speed-plot method [356]. The KH partial-wave analysis was focused on the intention to overcome the ambiguities due to incomplete experimental information by exploiting dispersion relations for fixed squared momentum transfer at small $t$, for fixed lab angles, and backward dispersion relations using $\bar{p}p\to\pi^{+}\pi^{-}$ data and the $\pi\pi$ phase shifts. Unitarity was imposed by adding an appropriate penalty term to the overall $\chi^{2}$ of the fit.
In the partial-wave analysis at CMB, real and imaginary parts of the invariant amplitudes were both parameterized by a sum of a term describing the Regge behavior at large energies and a series of “Born terms", representing $t$-channel Pomeron, $\rho$, and $f_{2}$-exchange and $u$-channel $N$ and $\Delta$ exchange. In the imaginary part, a diffractive Pomeron term was included and a suitable form representing two-pion exchange. Unitarity was enforced in the fits [357]. Ambiguities were resolved by forcing the amplitudes to be analytic in the Mandelstam plane along a few carefully chosen hyperbolas. In addition, some forward dispersion relations were used as part of the input. The resulting partial-wave amplitudes were described by coupled-channel Breit-Wigner amplitudes, whose pole positions and residues are presented and discussed in [358]. The results were updated at Baryon 1980 [359].
A long tradition of partial-wave analyses was initiated in 1969, when Arndt moved to the Virginia Polytechnic Institute (VPI) to collaborate with Roper [360]. Already at a very early stage, the group enabled external users to access the full data base and solution files through a Scattering Analysis Interactive Dial-in (SAID) computer system. In addition to their own solutions, SAID also encodes the Karlsruhe-Helsinki solution, the Carnegie-Mellon-Berkeley solution, and production partial waves. The system is an extremely useful tool to discuss experimental results and to plan new experiments.
@@
where $\rm T_{e}$ is the T-matrix of the energy-dependent analysis calculated at the given energy. The phase parameters $\delta$ and $\rho$ were expanded as linear functions around the given energy, with a slope (energy derivative) fixed by the energy-dependent solution. These single-energy results were then compared to the global fit in order to search for structure which could be missing. This procedure was iterated until a stable best-fit to the data was found. The solutions were constrained by dispersion relations. Milestones of their progress are documented in [361, 362, 363, 364, 365, 366, 367, 368, 369, 370, 371, 372, 373, 374, 375].
@@
A large number of groups, including SAID, have analyzed photo-induced reactions. Often, the intention of the work is a test of dynamical assumptions. Resonances are introduced with known properties, and just photocouplings are determined. In some cases, the agreement between different data is not really satisfactory. In the following we only discuss analyses that have been included by the Particle Data Group. An excellent summary of the methods applied by the different groups can be found in Ref. [14].
@@
The MAID group, led by Drechsel and Tiator at Mainz, developed isobar models for analyses of photo- and electroproduction of pions [124, 376, 377, 378, 379, 380], $\eta$ [381, 382, 383] and $\eta^{\prime}$ [384], two-pions [385], $\pi\eta$ [386], and $K\Lambda$ [387] production. MAID provides online programs [388] for real-time calculations of observables, amplitudes and partial waves (multipoles).
@@
The Kent State University group, led by Manley, contributed very significantly to the study of $N^{*}$ and $\Delta^{*}$ resonances and to the search for new states [40, 389, 390, 41, 391, 392, 393, 394]. In their latest publications, Hunt and Manley report a study of $\gamma p\to\eta p$, $\gamma n\to\eta n$ [392], and $\gamma p\to K^{+}\Lambda$ [393], and compare their partial-wave amplitudes with those obtained from other groups. The spread of results looks alarming. However, often the spread in masses, widths, and branching ratios is considerably smaller than one might suppose looking at the moments alone. In Ref. [394], Hunt and Manley present a fit to data on $\pi N\to\pi N$, $\to\pi\pi N$, $\to\eta N$, to $K\Lambda$, and to $\gamma N\to\pi N$, $\to\eta N$, and $\to K\Lambda$. A new $N(2200)7/2^{+}$ resonance with significant couplings to $\pi N$ and $K\Lambda$ is suggested. Masses, widths and decay properties of further 21 known $N^{*}$’s and twelve known $\Delta^{*}$’s are reported.
@@
Mosel at Gießen initiated coupled-channel analyses of pion and photoinduced reactions off nucleons in the late 70ies of the last century. The first fits included resonances with spin 1/2 and 3/2 only [395, 396, 397]. In [398], spin 5/2 was introduced. The further fits included eleven $N^{*}$ and nine $\Delta^{*}$ resonances with spins $J\leq 5/2$ and reported masses, width and branching ratios for decays into $\pi N$ [399], $K\Lambda$ [400], $\eta N$ [401, 402], and $K\Sigma$ [403]. No errors were evaluated. In Ref. [404], the emphasis of the analysis is to study the isobars in the $N(1440)\to N\pi\pi$ decays.
@@
The Bonn-Gatchina (BnGa) group was initiated in 2002 by one of the authors (E.K.). The group performed coupled-channel analyses of a very large data set on pion and photo-induced reactions and discovered several new resonances or improved the evidence for their existence. The method is described in [405, 406, 407] and results are presented in Refs. [160, 165, 166, 167, 168, 179, 185, 408, 409, 410, 411, 412, 413, 414, 415, 416, 417, 418, 419, 420, 421, 422, 423, 424, 425, 426, 427]. The distinctive features of the BnGa analysis are the use of event-based likelihood fits to three-body final states and the rapid convergence of the fitting procedure. This aim is achieved by implementing derivatives of fit quantities and a multilevel structure for fitting parameters. The speed of the fit allows the group to systematically search for new resonances in all partial waves. Systematic errors are evaluated by adding further resonances at high masses and by changing the weights with which data sets are included in the fits. The group found that including or excluding high-mass resonances provides the largest contribution to the final uncertainties. The data base and fit curves can be found on the web page [428].
@@
The ANL-Osaka group deveolped a dynamical coupled-channel model for the combined analysis of pion- and photon-induced reactions. The progress is reported in a series of publications [429, 430, 431, 432, 433, 434, 435, 436, 437, 438, 439, 440, 441, 442, 443, 444, 445]. In the latest works, a very large data set on pion and photo-induced reactions is used to determine properties of ten $N^{*}$ and nine $\Delta^{*}$ resonances. The approach concentrates on a rigorous definition of the scattering and production amplitudes, but at the cost of a large amount of computing time which limits the feasability of systematic studies.
@@
The Jülich-Bonn (JüBo) group developed a dynamical coupled-channel model for the combined analysis of a very large data set on pion- and photo-induced reactions, referred to as the Jülich-Bonn-Washington model [446, 447, 448]. First, the analytic properties of a resonance were discussed [449], the dependence of the pole properties on the background amplitudes [450], and the phase and pole structure of $N(1535)1/2^{-}$ [451]. In a study of data on $I_{s}$, $I_{c}$ for the reaction $\gamma p\to\pi^{0}\eta p$ [452], the $\Delta(1700)3/2^{-}$ was interpreted as a dynamically generated resonance. In a series of publications, an increasing data base was included in a coupled-channel fit [453, 454, 455, 456, 457, 458] resulting in a full spectrum of $N^{*}$ and $\Delta^{*}$ resonances with their decay properties. Figures representing the fit results including a display of the data can be found in Ref. [459]. The dynamical coupled-channel approaches applied by the ANL-Osaka and Jülich-Bonn groups are reviewed in Ref. [460].
@@
A simple method to extract pole parameters from the data, called the Laurent-Pietarinen (L + P) method, has been developed by Švarc and collaborators. Data, mostly single channels, were fitted with a function containing poles, branch points and cuts [461, 462, 463, 464, 465, 466, 283, 167, 467, 468, 469, 470, 471, 472, 473]. The results are pole positions, but no branching ratios or transition residues have been derived.
@@
Figure 2.1: Comparison of PWA results on pole positions ($M$ and $\Gamma$) of $N^{*}$ (left) and $\Delta^{*}$ (right) resonances, including JüBo [458], BnGa [474, 475], Kent [394], SAID [365], and the RPP 2024 [3]. Graphics credit: Ulrike Thoma.
2.2.2   Hyperon resonances
Most hyperon resonances were discovered in the 70s of the last century in single-channel analyses. The results obtained so far were summarized by Gopal at Baryon 1980, the 4th International Conference on Baryon Resonances, held in Toronto, Canada, in July 1980 [476]. Then, the field had a standstill. Manley and his group at KSU collected a large set of data on $K^{-}p$ interactions and performed a coupled-channel analysis [477, 478]. The KSU partial-wave amplitudes were also fitted by the JPAC group in a coupled-channel fit [479]. The Osaka-ANL group analyzed the same data in a dynamical coupled-channel approach [480, 481]. The BnGa group added further published data and performed a coupled-channel analysis [47, 482].
Figure 2.2: (color online) Light and strange baryon spectrum from the PDG [3]. The columns correspond to different $J^{P}$ and the colors to isospin and hypercharge. The different font weights represent four-, three-, two- and one-star resonances.
Table 2.1: $\chi^{2}$ per data point for the consistency of the results on the pole positions of 4* resonances from different PWAs with the RPP mean values within the estimated ranges. BnGa = Bonn-Gatchina, JüBo = Jülich-Bonn, CM = Carnegie-Mellon, PiA = Pittsburgh-Argonne, KH = Karlsruhe-Helsinki Collaboration. In most analyses, $N_{\rm data}=28$ or 30.
BnGa [475]: 0.66 JüBo [458]: 3.96 BnGa [483]: 0.55
L+P [463]: 0.47 CM [359]: 0.78 Kent [394]: 1.28
JüBo [456]: 4.06 Gi [402]: 8.11 BnGa [419]: 0.77
SAID [365]: 1.58 PiA [484]: 4.36 KH [27]: 1.18

### v4

@@ 2.1  Data
Electroproduction: Meson photoproduction has become an essential tool in the search for new excited light-quark baryon states. As discussed in the previous section, many new excited states have been discovered thanks to high-precision photoproduction data in different final states Anisovich:2017bsk , and are now included in recent editions of the Review of Particle Physics (RPP) ParticleDataGroup:2022pth . The experimental exploration of resonance transition form factors spans over 60 years, with many review articles describing this history. Here we refer to some recent ones Aznauryan:2011qj , Stoler:1993yk , Burkert:2004sk , Aznauryan:2012ba . A review of recent electroproduction experiments in hadron physics and their interpretation within modern approaches to strong interaction physics can be found in Ref. Proceedings:2020fyd . The exploration of the internal structure of excited states and the effective degrees of freedom contributing to $s$-channel resonance excitation requires the use of electron beams, where the virtuality ($Q^{2}$) of the exchanged photon can be varied to penetrate through the peripheral meson cloud and probe the quark core and its spatial structure. Electroproduction can thus reveal whether a resonance is generated through short-distance photon interaction with the small quark core or through interaction with a more extended hadronic system. Electroproduction of final states with pseudoscalar mesons (e.g., $N\pi$, $p\eta$, $K\Lambda$, $p\pi^{-}\pi^{+}$) has been employed at Jefferson Laboratory mainly with the CEBAF Large Acceptance Spectrometer (CLAS) operating at an instantaneous luminosity of $10^{34}\,\mathrm{sec}^{-1}\mathrm{cm}^{-2}$. In Hall A and Hall C, pairs of individual well-shielded focusing magnetic spectrometers are employed with more specialized aims and limited acceptance, but operating at much higher luminosity. This experimental program has led to new insights into the scale dependence of effective degrees of freedom, e.g., meson-baryon, constituent quarks, and dressed quark contributions. Data on meson electroproduction can be found in Refs. CLAS:2023mfc , Mokeev:2023zhq , CLAS:2022iqy , CLAS:2022kta , CLAS:2021cvy , Mokeev:2020hhu , CLAS:2021jhm , CLAS:2021ovm , CLAS:2019cpp , CLAS:2019uzc , CLAS:2018fon , Aznauryan:2018okk , Aznauryan:2017nkz , Aznauryan:2016wwm , CLAS:2017rgp , CLAS:2017fja , CLAS:2017jjr , CLAS:2016tqs , CLAS:2016cks , CLAS:2016dlz , CLAS:2016qxj , CLAS:2016ikd , Kim:2015pkf , Skorodumina:2015ccu , Mokeev:2015lda , CLAS:2014fml , CLAS:2014jpc , CLAS:2014udv , Aznauryan:2015zta , Aznauryan:2014xea , CLAS:2012qga , Aznauryan:2012ba , CLAS:2012wxw , CLAS:2012cna , Aznauryan:2012ec , Aznauryan:2011qj , CLAS:2009ces , CLAS:2009sbn , CLAS:2008roe , CLAS:2008wls , CLAS:2008agj , CLAS:2007jvi , CLAS:2008ihz , CLAS:2008rpm , CLAS:2008cms , CLAS:2007bvs , CLAS:2006sjw , CLAS:2006ogr , CLAS:2006ezq , CLAS:2005oqk , Aznauryan:2005tp , CLAS:2005vxa , CLAS:2005nkx , CLAS:2004cri , Aznauryan:2004jd , CLAS:2004ncx , Burkert:2004sk , CLAS:2002xbv , CLAS:2002rxi , CLAS:2003hro , CLAS:2002zlc , CLAS:2001cbm , CLAS:2000mbw , Breuker:1977vy , Breuker:1982nw , Beck:1974wd , Siddle:1971ug , Alder:1975tv , Brasse:1975bg , Alder:1975na , Alder:1975xt , Brasse:1976bf , Brasse:1977as , Haidan:1979yqa .
@@ 2.2  Analyses
The main KH results can be found in Hohler:1979yr , with an update of the results presented in Koch:1980ae . A comprehensive description of methods and results can be found in the handbook Hohler:1984ux published in the Landold-Börnstein series. Höhler presented updated resonance masses and widths using the so-called speed-plot method Hohler:1992ru . The KH partial-wave analysis was focused on the intention to overcome the ambiguities due to incomplete experimental information by exploiting dispersion relations for fixed squared momentum transfer at small $t$, for fixed lab angles, and backward dispersion relations using $\bar{p}p\to\pi^{+}\pi^{-}$ data and the $\pi\pi$ phase shifts. Unitarity was imposed by adding an appropriate penalty term to the overall $\chi^{2}$ of the fit.
In the partial-wave analysis at CMB, real and imaginary parts of the invariant amplitudes were both parameterized by a sum of a term describing the Regge behavior at large energies and a series of “Born terms", representing $t$-channel Pomeron, $\rho$, and $f_{2}$-exchange and $u$-channel NN and $\Delta$ exchange. In the imaginary part, a diffractive Pomeron term was included and a suitable form representing two-pion exchange. Unitarity was enforced in the fits Cutkosky:1979fy . Ambiguities were resolved by forcing the amplitudes to be analytic in the Mandelstam plane along a few carefully chosen hyperbolas. In addition, some forward dispersion relations were used as part of the input. The resulting partial-wave amplitudes were described by coupled-channel Breit-Wigner amplitudes, whose pole positions and residues are presented and discussed in Cutkosky:1979zv . The results were updated at Baryon 1980 Cutkosky:1980rh .
A long tradition of partial-wave analyses was initiated in 1969, when Arndt moved to the Virginia Polytechnic Institute (VPI) to collaborate with Roper Arndt:1970wr . Already at a very early stage, the group enabled external users to access the full data base and solution files through a Scattering Analysis Interactive Dial-in (SAID) computer system. In addition to their own solutions, SAID also encodes the Karlsruhe-Helsinki solution, the Carnegie-Mellon-Berkeley solution, and production partial waves. The system is an extremely useful tool to discuss experimental results and to plan new experiments.
@@
where $\rm T_{e}$ is the T-matrix of the energy-dependent analysis calculated at the given energy. The phase parameters $\delta$ and $\rho$ were expanded as linear functions around the given energy, with a slope (energy derivative) fixed by the energy-dependent solution. These single-energy results were then compared to the global fit in order to search for structure which could be missing. This procedure was iterated until a stable best-fit to the data was found. The solutions were constrained by dispersion relations. Milestones of their progress are documented in Arndt:1985vj , Arndt:1990bp , Arndt:1995bj , Arndt:2003if , Arndt:2006bf , Workman:2008iv , Workman:2011vb , Workman:2012jf , Workman:2012hx , Briscoe:2012ni , Workman:2016ysf , Briscoe:2020qat , Briscoe:2021siu , Strakovsky:2022tvu , Briscoe:2023gmb .
@@
A large number of groups, including SAID, have analyzed photo-induced reactions. Often, the intention of the work is a test of dynamical assumptions. Resonances are introduced with known properties, and just photocouplings are determined. In some cases, the agreement between different data is not really satisfactory. In the following we only discuss analyses that have been included by the Particle Data Group. An excellent summary of the methods applied by the different groups can be found in Ref. Thiel:2022xtb .
@@
The MAID group, led by Drechsel and Tiator at Mainz, developed isobar models for analyses of photo- and electroproduction of pions A2:2015mhs , Drechsel:1998hk , Kamalov:2000en , Tiator:2008kd , Tiator:2010rp , Stajner:2017fmh , $\eta$ A2atMAMI:2014rwc , Drechsel:2007if , Chiang:2001as and $\eta^{\prime}$ Tiator:2018heh , two-pions Fix:2005if , $\pi\eta$ Fix:2010nv , and $K\Lambda$ Mart:1999ed production. MAID provides online programs MAIDkph for real-time calculations of observables, amplitudes and partial waves (multipoles).
@@
The Kent State University group, led by Manley, contributed very significantly to the study of $N^{*}$ and $\Delta^{*}$ resonances and to the search for new states Manley:1984zs , Manley:1984ih , Manley:1984jz , Manley:1992yb , Shrestha:2012va , Hunt:2018tvt , Hunt:2018mrt , Hunt:2018wqz . In their latest publications, Hunt and Manley report a study of $\gamma p\to\eta p$, $\gamma n\to\eta n$ Hunt:2018tvt , and $\gamma p\to K^{+}\Lambda$ Hunt:2018mrt , and compare their partial-wave amplitudes with those obtained from other groups. The spread of results looks alarming. However, often the spread in masses, widths, and branching ratios is considerably smaller than one might suppose looking at the moments alone. In Ref. Hunt:2018wqz , Hunt and Manley present a fit to data on $\pi N\to\pi N$, $\to\pi\pi N$, $\to\eta N$, to $K\Lambda$, and to $\gamma N\to\pi N$, $\to\eta N$, and $\to K\Lambda$. A new $N(2200)7/2^{+}$ resonance with significant couplings to $\pi N$ and $K\Lambda$ is suggested. Masses, widths and decay properties of further 21 known $N^{*}$’s and twelve known $\Delta^{*}$’s are reported.
@@
Mosel at Gießen initiated coupled-channel analyses of pion and photoinduced reactions off nucleons in the late 70ies of the last century. The first fits included resonances with spin 1/2 and 3/2 only Penner:2001fu , Penner:2002ma , Penner:2002md . In Shklyar:2004dy , spin 5/2 was introduced. The further fits included eleven $N^{*}$ and nine $\Delta^{*}$ resonances with spins $J\leq 5/2$ and reported masses, width and branching ratios for decays into $\pi N$ Shklyar:2004ba , $K\Lambda$ Shklyar:2005xg , $\eta N$ Shklyar:2006xw , Shklyar:2012js , and $K\Sigma$ Cao:2013psa . No errors were evaluated. In Ref. Shklyar:2014kra , the emphasis of the analysis is to study the isobars in the $N(1440)\to N\pi\pi$ decays.
@@
The Bonn-Gatchina (BnGa) group was initiated in 2002 by one of the authors (E.K.). The group performed coupled-channel analyses of a very large data set on pion and photo-induced reactions and discovered several new resonances or improved the evidence for their existence. The method is described in Anisovich:2004zz , Anisovich:2006bc , Burkert:2022bqo and results are presented in Refs. Anisovich:2018yoo , Anisovich:2017pox , Anisovich:2015gia , Anisovich:2017ygb , Anisovich:2017xqg , Anisovich:2014yza , Anisovich:2013jya , Sarantsev:2005tg , Anisovich:2005tf , Klempt:2006sa , Nikonov:2007br , Anisovich:2007bq , Anisovich:2008wd , Anisovich:2010mks , Anisovich:2010an , Anisovich:2011ye , Anisovich:2011sv , Anisovich:2011su , Anisovich:2011fc , Anisovich:2012ct , Anisovich:2013sva , Anisovich:2013tij , Anisovich:2013vpa , Anisovich:2015tla , Denisenko:2016ugz , Anisovich:2016vzt , Anisovich:2017afs . The distinctive features of the BnGa analysis are the use of event-based likelihood fits to three-body final states and the rapid convergence of the fitting procedure. This aim is achieved by implementing derivatives of fit quantities and a multilevel structure for fitting parameters. The speed of the fit allows the group to systematically search for new resonances in all partial waves. Systematic errors are evaluated by adding further resonances at high masses and by changing the weights with which data sets are included in the fits. The group found that including or excluding high-mass resonances provides the largest contribution to the final uncertainties. The data base and fit curves can be found on the web page pwahiskp .
@@
The ANL-Osaka group deveolped a dynamical coupled-channel model for the combined analysis of pion- and photon-induced reactions. The progress is reported in a series of publications Sato:1996gk , Sato:2000jf , Julia-Diaz:2006ios , Julia-Diaz:2007qtz , Matsuyama:2006rp , Suzuki:2008rp , Kamano:2008gr , Julia-Diaz:2009dnz , Suzuki:2009nj , Kamano:2009im , Sandorfi:2010uv , Suzuki:2010yn , Kamano:2010ud , Kamano:2013iva , Nakamura:2015rta , Kamano:2016bgm , Kamano:2019gtm . In the latest works, a very large data set on pion and photo-induced reactions is used to determine properties of ten $N^{*}$ and nine $\Delta^{*}$ resonances. The approach concentrates on a rigorous definition of the scattering and production amplitudes, but at the cost of a large amount of computing time which limits the feasability of systematic studies.
@@
The Jülich-Bonn (JüBo) group developed a dynamical coupled-channel model for the combined analysis of a very large data set on pion- and photo-induced reactions, referred to as the Jülich-Bonn-Washington model Mai:2021vsw , Mai:2021aui , Mai:2023cbp . The analytic properties of a resonance were discussed Doring:2009yv and the dependence of the pole properties on the background amplitudes Doring:2009bi . The phase and pole structure of $N(1535)1/2^{-}$ by Döring and Nakayama Doring:2009uc . In a study of data on $I_{s}$, $I_{c}$ for the reaction $\gamma p\to\pi^{0}\eta p$ CBELSATAPS:2015tyg , prominent contributions due to $\Delta(1700)3/2^{-}$ formation was found, interpreted as a dynamically generated resonance in Ref. Doring:2005bx .
In a series of publications, the JüBo group included an increasing data base in a coupled-channel fit Doring:2010ap , Ronchen:2012eg , Ronchen:2014cna , Ronchen:2015vfa , Ronchen:2018ury , Ronchen:2022hqk resulting in a full spectrum of $N^{*}$ and $\Delta^{*}$ resonances with their decay properties. Figures representing the fit results including a display of the data can be found in Ref. collaborations.fz-juelich . The dynamical coupled-channel approaches applied by the ANL-Osaka and Jülich-Bonn groups are reviewed in Ref. Doring:2025sgb .
@@
A simple method to extract pole parameters from the data, called the Laurent-Pietarinen (L + P) method, has been developed by Švarc and collaborators. Data, mostly single channels, were fitted with a function containing poles, branch points and cuts Svarc:2012pt , Svarc:2013laa , Svarc:2014zja , Svarc:2014sqa , Svarc:2014aga , Svarc:2015usk , Anisovich:2017bsk , Anisovich:2017ygb , Svarc:2018eol , Svarc:2018aay , Svarc:2020waq , Svarc:2020cic , Svarc:2021zkb , Svarc:2021gcs , Svarc:2022buh . The results are pole positions, but no branching ratios or transition residues have been derived.
In photoproduction, the nucleon and the photon spin can be parallel or opposite, the nucleon spin can flip or not. Thus there are four independent amplitudes describing the process. In single pseudoscalar meson photoproduction, there are sixteen linearly independent measurable quantities; the four complex amplitudes require at least eight independent measurements. In energy-dependent analyses, these four amplitudes are constrained by a model. In the recent years, progress has been made in energy-independent reconstructions of amplitudes in truncated partial-wave analyses in which the number of required measurements is reduced to five Wunderlich:2013iga , Workman:2016irf . The method has been applied to single-pion photoproduction Fix:2022shn , to $\gamma p\to\Lambda K^{+}$ Svarc:2021gcs , and to $\gamma p\to p\eta$ Svarc:2020cic , Kroenert:2023ovd . In the future, the constructed amplitudes could be used to extract the pole position of resonances in the L + P formalism.
@@
Figure 2.1: Comparison of PWA results on pole positions ($M$ and $\Gamma$) of $N^{*}$ (left) and $\Delta^{*}$ (right) resonances, including JüBo Ronchen:2022hqk , BnGa CLAS:2024iir , Sarantsev:2025lik , Kent Hunt:2018wqz , SAID Arndt:2006bf , and the RPP 2024 ParticleDataGroup:2024cfk . Graphics credit: Ulrike Thoma.
Hyperon resonances: Most hyperon resonances were discovered in the 70s of the last century in single-channel analyses. The results obtained so far were summarized by Gopal at Baryon 1980, the 4th International Conference on Baryon Resonances, held in Toronto, Canada, in July 1980 Gopal:1980ur . Then, the field had a standstill. Manley and his group at KSU collected a large set of data on $K^{-}p$ interactions and performed a coupled-channel analysis Zhang:2013cua , Zhang:2013sva . The KSU partial-wave amplitudes were also fitted by the JPAC group in a coupled-channel fit Fernandez-Ramirez:2015tfa . The Osaka-ANL group analyzed the same data in a dynamical coupled-channel approach Kamano:2014zba , Kamano:2015hxa . The BnGa group added further published data and performed a coupled-channel analysis Matveev:2019igl , Sarantsev:2019xxm .
Figure 2.2: (color online) Light and strange baryon spectrum from the PDG ParticleDataGroup:2024cfk . The columns correspond to different $J^{P}$ and the colors to isospin and hypercharge. The different font weights represent four-, three-, two- and one-star resonances.
Table 2.1: $\chi^{2}$ per data point for the consistency of the results on the pole positions of 4* resonances from different PWAs with the RPP mean values within the estimated ranges. BnGa = Bonn-Gatchina, JüBo = Jülich-Bonn, Gi = Gießen, CM = Carnegie-Mellon, PiA = Pittsburgh-Argonne, KH = Karlsruhe-Helsinki. In most analyses, $N_{\rm data}=28$ or 30.
BnGa Sarantsev:2025lik : 0.66 JüBo Ronchen:2022hqk : 3.96 BnGa CBELSATAPS:2015kka : 0.55
L+P Svarc:2014zja : 0.47 CM Cutkosky:1980rh : 0.78 Kent Hunt:2018wqz : 1.28
JüBo Ronchen:2015vfa : 4.06 Gi Shklyar:2012js : 8.11 BnGa Anisovich:2011fc : 0.77
SAID Arndt:2006bf : 1.58 PiA Vrana:1999nt : 4.36 KH Hohler:1993lbk : 1.18